# Decentralized and Uncoordinated Learning of Stable Matchings: A Game-Theoretic Approach

### v1

@@ Abstract
We consider the problem of learning stable matchings in a fully decentralized and uncoordinated manner, where decentralized means that players make decisions individually without the influence of a central platform, and uncoordinated means that players do not need to synchronize their decisions using pre-specified rules. In this problem, there are $n$ men and $n$ women, each having preference over the other side. It is assumed that women know their preferences over men, but men are not aware of their preferences over women, and they only learn them if they propose and successfully get matched to women. A matching is called stable if no man and woman prefer each other over their current matches. When all the preferences are known a priori, the celebrated Deferred-Acceptance (DA) algorithm proposed by Gale and Shapley provides a decentralized and uncoordinated algorithm to obtain a stable matching. However, when the preferences are unknown, developing such an algorithm faces major challenges due to a lack of coordination. In this work, we achieve this goal by making a connection between stable matchings and learning Nash equilibria (NE) in noncooperative games. First, we provide a complete information game formulation for the stable matching problem with known preferences such that its set of pure NE coincides with the set of stable matchings, while its mixed NE with full support can be rounded in a decentralized manner to a stable matching. Relying on such a game-theoretic formulation, we show that for hierarchical structured markets, adopting the exponential weight (EXP) learning algorithm for the stable matching game achieves logarithmic regret with polynomial dependence on the number of players, thus answering a question posed in previous literature. Moreover, we show that the same EXP learning algorithm converges locally and exponentially fast to a stable matching in general matching markets. We complement this result by introducing another decentralized and uncoordinated learning algorithm that globally converges to a stable matching with arbitrarily high probability, leveraging the weak acyclicity property of the stable matching game. Since the global convergence of this algorithm can be slow in general matching markets, we investigate conditions under which stable matchings can be learned more quickly. In particular, for general matching markets, we show that if men receive stronger feedback when their proposals are rejected, it is possible to drive the market more rapidly toward an approximate stable matching. Our proposed game-theoretic framework bridges the discrete problem of learning stable matchings with the problem of learning NE in continuous-action games.
@@
Stable matchings constitute one of the fundamental problems in computer science, economics, and engineering that have received considerable attention in the past decades. Stable matchings provide a desirable notion of stability in two-sided matching markets where the agents on each side of the market have preferences over the other side. For instance, in college admission, the applicants have different preferences over colleges and vice versa, and the goal is to match the applicants to the colleges so that no applicant-college pair would prefer to break their current matches and instead get matched to each other. Similar situations arise in other applications, such as kidney exchange programs, assigning jobs to workers, matching in online dating platforms, and scheduling jobs on heterogeneous machines.
@@
It is known that when all the preferences are known, stable matchings always exist and a simple decentralized and uncoordinated Deferred-Acceptance (DA) algorithm, first proposed by [1], can find one such stable matches in polynomial number of iterations. Here, by a decentralized algorithm we refer to an algorithm in which players make decisions individually without the influence of a central platform, and uncoordinated means that players do not need to synchronize their decisions using pre-specified rules. While the DA algorithm provides a satisfactory solution to many practical applications, there are scenarios where either the information structure of the problem hinders the implementation of the DA algorithm or the stable matching achieved by the DA algorithm is inferior in terms of social optimality, especially when there exist many feasible stable matchings in the market. For instance, there are many recent emergence of online matching markets such as online labor markets (e.g., TaskRabbit, Upwork), online dating markets (e.g., Tinder, Match.com), and online crowdsourcing platforms (e.g., Amazon mechanical turk), where the users do not know their preferences a priori, and can repeatedly interact with the market to improve their matching quality [2]. The more a pair on both sides of the market gets matched, the more certain they will be about their preferences. Therefore, a question of interest is how the agents should interact with the market such that in the absence of any coordination, they will learn their preferences as quickly as possible, and their collective behavior results in a stable matching.
@@
In this work, we consider a two-sided matching market posed as a marriage problem, where we refer to the agents of one side as men and to the agents of the other side as women. We assume both men and women have preferences on the other side and that women are aware of their ordinal preferences about men. However, men do not know their preferences about women and only learn those preferences if they propose and get matched, in which case the matched men receive a noisy estimate of their preferences. We assume that men/women cannot observe any other information (e.g., who is rejected/accepted), nor can they coordinate in any sense. The goal for men is to follow a decentralized and uncoordinated proposal strategy such that the entire market converges to a stable matching over time.
@@
One of the major challenges for learning a stable matching in markets with unknown preferences is the collisions. More precisely, when multiple men propose to the same woman, only one of them gets matched and receives useful information; all others get rejected and receive no information regarding their preferences for that woman. Therefore, resolving such collisions without coordination is a major issue. Moreover, even if a man gets matched to a woman, he receives a noisy utility drawn from an unknown distribution that could differ from his true preference. Thus, a collision-avoidance process should be repeated many times until men can accurately estimate their true preferences. One of our goals in this work is to provide decentralized and uncoordinated algorithms that can effectively learn the true preferences while incurring a minimum number of collisions. Moreover, our work provides an alternative approach for measuring the closeness of the generated dynamics to any stable matching in general matching markets and not necessarily to a specific stable matching in structured markets, which has been the main focus in the past literature. To that end, we make a connection between learning stable matchings in general markets with unknown preferences and learning Nash equilibrium (NE) in noncooperative games and leverage this connection to devise decentralized and uncoordinated algorithms for learning stable matchings.
-A Related Work
@@
Stable matching was first introduced by [1] as a model for college admissions and to study the stability of marriages. Since then, there has been a tremendous effort on generalization and extension of stable matchings to various market settings [3]. In their seminal work, Gale and Shapley provided a simple DA algorithm in which men propose to their most preferred women, and the women reject all the received proposals except the most preferred one. They show that such a decentralized algorithm converges to a stable matching in at most $O(n^{2})$ steps, where $n$ is the total number of men/women in the market. Unfortunately, when the agents’ preferences are unknown, the DA algorithm is no longer guaranteed to converge to a stable matching without extra coordination. More precisely, if men fail to propose to their highest preferred women (due to a lack of knowledge of their true preferences), their subsequent proposals may become unpredictable. One reason is that if a man $m$ is rejected by a woman $w$ at a time instance $t$, that does not mean that $m$ should stop proposing to $w$ because it could be the case that another man mistakenly proposed to $w$ at time $t$, hence causing the rejection of $m$ by $w$. Therefore, in order to extend the DA algorithm to the case of unknown preferences, one requires some sort of coordination among the agents to avoid uninformative acceptance/rejection patterns. To address this issue, the earlier works [4, 2, 5, 6] have proposed various coordination-based DA algorithms, where the main idea is to develop a phase-dependent process where in some phases the agents mostly explore, and in other phases, they implement the DA algorithm based on the current estimates of their unknown preferences.
@@
It is known that stable matchings with known preferences can be characterized using extreme points of a fractional polytope that is totally dual integral [7, 8]. Moreover, stable matchings exhibit other nice properties that allow one to characterize their geometry using so-called rotations [9]. It was shown in [10] that uncoordinated random better response dynamics converge to a stable matching with probability one. Subsequently, [11] provided an exponential lower bound for the worst-case convergence time of the uncoordinated random better/best response dynamics to a stable matching. However, all these results are for the case when the market preferences are fully known so that the agents can compute their best/better responses and update their decisions accordingly.
@@
More recently, there has been a huge interest in matching markets with unknown preferences. However, depending on the information structure and the feedback received by the agents, one would expect a wide range of performance guarantees, which also depends on the type of algorithms followed by the agents (e.g., centralized/decentralized or coordinated/uncoordinated). The work [12] devised a randomized polynomial-time centralized algorithm for finding a stable matching with unknown deterministic preferences, where each time, the algorithm proposes a matching and receives feedback in terms of a blocking pair. [^1] A unmatched pair of man-woman $(m,w)$ is called blocking if both $m$ and $w$ prefer each other to their matched partner. The work [13] considers learning a specific stable matching with unknown stochastic preferences by adopting a suitable notion of stable regret that measures the number of time instances that men propose to women other than their stable pair in that stable matching. [2] considered the problem of learning stable matchings in hierarchically structured markers [^2] In fact, the hierarchical condition (a.k.a. sequential preference condition [14]) that we consider in Section III is weaker than the $\alpha$-reducible condition considered in [2], and our results hold for broader matching markets including the ones considered in [2]. in which each submarket has a fixed pair (i.e., a pair of men-women that prefer each other the most in that submarket), and developed a phase-coordinated decentralized algorithm that achieves logarithmic regret with respect to the market’s unique stable matching. However, the regret bound in [2] depends exponentially on other parameters such as the number of men/women. Furthermore, [15] considered learning stable matchings in an uncoordinated and decentralized fashion. However, they require stronger assumptions on the information feedback (e.g., a player observes the actions of other players in the previous round) and use a different performance metric than the one we consider in this work. We refer to [16] and [17] for other learning algorithms in two-sided matching markets with different performance guarantees. While all these works essentially address the same problem of learning a stable matching with unknown preferences, their algorithms differ substantially due to the information/feedback structure, the type of performance guarantee, and the amount of allowable coordination among the agents.
-B Contributions and Organization
@@
We consider the problem of learning stable matchings in matching markets, where men do not know their preferences and want to achieve a stable matching in a decentralized and uncoordinated fashion. In particular, we consider the weakest type of feedback that men can receive: a man gets to observe a noisy version of his true preference only if his proposal gets accepted and receives no information otherwise. Our main objective is to provide a novel and principled approach to designing decentralized and uncoordinated learning algorithms for stable matchings through the lens of NE learning in noncooperative games. Our contributions can be summarized as follows:
@@
  • We first provide a complete information continuous-action game formulation for the stable matching problem and show that its set of pure NE coincides with the set of stable matchings while its mixed NE points can be rounded in a decentralized way to a stable matching. This connection provides an alternative approach for measuring the closeness of the market dynamics to stable matchings through the lens of NE computation in games, hence extending the conventional notion of regret with respect to a specific stable matching to any arbitrary stable matching in general markets.
  • Leveraging this game-theoretic formulation, we provide a simple decentralized and uncoordinated algorithm EXP that converges globally to a stable matching in hierarchical markets and achieves logarithmic regret with only polynomial dependence on the number of players, hence improving the existing bounds in the past literature.
@@
  • We then show that the EXP algorithm converges locally exponentially fast to a stable matching in general matching markets. In particular, we show that if EXP converges with positive probability, it converge globally to a stable matching. [^3] Perhaps fast local convergence is the best one can hope for general markets because, for general matching markets, there are strong exponential lower bounds for the number of iterates of the better response dynamics to converge globally to a stable matching.
@@
  • We also provide an alternative decentralized and uncoordinated algorithm that globally converges to a stable matching in general markets with arbitrarily high probability (albeit at a low convergence rate), thanks to the weakly acyclic property of the stable matching game.
  • Finally, we address an information feedback design problem by seeking conditions in general matching markets that allow players to learn a stable matching quickly. In this regard, we leverage our game-theoretic framework to provide one such information feedback by establishing a connection to monotone games. Our results offer new insights into designing payoff functions that enable faster decentralized and uncoordinated algorithms for driving matching markets to their stable points.
@@
The paper is organized as follows. In Section I, we formally introduce the problem. In Section II, we provide a complete information game-theoretic formulation for the stable matching problem and prove several characterization results for its set of NE points. In Section III, we provide a decentralized and uncoordinated algorithm (EXP) that achieves logarithmic regret for hierarchical markets with polynomial dependence on the number of players. In Section IV, we show that EXP converges locally exponentially fast to a stable matching in general matching markets. In Section V, we provide an alternative decentralized and uncoordinated algorithm that globally converges to a stable matching in general matching markets with arbitrarily high probability. In Section VI, we extend our results to matching markets with stronger feedback and establish a connection between fast global learning of a stable matching and learning a NE in monotone games. We conclude the paper by identifying some future research directions in Section VII.
I Problem Formulation
Stable Matchings with Known Preferences
@@ I Problem Formulation
Here, we first introduce the stable matching problem with known preferences [1]. In this problem, there are a set $M$ of men and a set $W$ of women such that $|M|=|W|=n$ (we use the term “agents” to refer to either men or women). [^4] By introducing dummy agents with appropriate preferences, without loss of generality we may assume $|M|=|W|$ [7]. Each man $m\in M$ has a cardinal preference for each women $w\in W$, denoted by $\mu_{mw}\in(0,1]$, such if $\mu_{mw}>\mu_{mw^{\prime}}$, it means that $m$ prefers $w$ over $w^{\prime}$. Moreover, each woman $w$ has an ordinal preference over the men, and we write $m>_{w}m^{\prime}$ if woman $w$ strictly prefers $m$ over $m^{\prime}$. In this work we assume that no agent has ties in their preferences and we define
@@
Stable Matchings with Unknown Preferences
@@
II A Game-Theoretic Formulation and Preliminary Results
In this section, we provide a complete information noncooperative game formulation for the stable matching problem with known preferences whose set of pure Nash equilibrium (NE) points coincides with the set of stable matchings. Later, we will show how to leverage this formulation to extend our results for learning stable matchings with unknown preferences. Such a formulation has three main advantages: i) it reduces the learning task in matching markets to learning NE in continuous-action concave games, ii) it captures the selfish behavior of men (players) and the feedback they receive through their payoff functions, and iii) it simplifies the combinatorial structure inherent in learning stable matchings to learning NE in continuous-action games. While analyzing the agents’ decisions directly on the discrete space of matchings seems complicated, which is why some earlier works need to restrict attention to special market structures [2], our game-theoretic reduction simplifies this task to some extent, thanks to the extensive literature on NE learning in concave games.

### v2

@@ Abstract
We consider the problem of learning stable matchings with unknown preferences in a decentralized and uncoordinated manner, where “decentralized” means that players make decisions individually without the influence of a central platform, and “uncoordinated” means that players do not need to synchronize their decisions using pre-specified rules. First, we provide a game formulation for this problem with known preferences, where the set of pure Nash equilibria (NE) coincides with the set of stable matchings, and mixed NE can be rounded to a stable matching. Then, we show that for hierarchical markets, applying the exponential weight (EXP) learning algorithm to the stable matching game achieves logarithmic regret in a fully decentralized and uncoordinated fashion. Moreover, we show that EXP converges locally and exponentially fast to a stable matching in general markets. We also introduce another decentralized and uncoordinated learning algorithm that globally converges to a stable matching with arbitrarily high probability. Finally, we provide stronger feedback conditions under which it is possible to drive the market faster toward an approximate stable matching. Our proposed game-theoretic framework bridges the discrete problem of learning stable matchings with the problem of learning NE in continuous-action games.
@@ I Introduction
Learning stable matchings is one of the fundamental problems in computer science, economics, and engineering that has received considerable attention over the past decades. Stable matchings provide a desirable notion of stability in two-sided matching markets where agents on each side of the market have preferences over the other side. A matching is called stable if no two agents prefer each other over their current matches. For instance, in college admissions, applicants have different preferences for colleges and vice versa. The goal is to match applicants to colleges in such a way that no applicant-college pair would prefer to break their current matches and instead be matched to each other. Similar situations arise in other applications, such as kidney exchange programs, job assignment to workers, matching in online dating platforms, and scheduling jobs on heterogeneous machines.
@@
It is known that when all preferences are known, stable matchings always exist, and a simple decentralized and uncoordinated Deferred-Acceptance (DA) algorithm, first proposed by [1], can find such stable matches in a polynomial number of iterations. Here, a decentralized algorithm refers to one in which players make decisions individually without the influence of a central platform, and uncoordinated means that players do not need to synchronize their decisions using pre-specified rules. However, when preferences are unknown, developing such an algorithm faces major challenges due to the lack of coordination. While the DA algorithm provides a satisfactory solution for many practical applications, there are scenarios where either the information structure of the problem hinders the implementation of the DA algorithm or the stable matching achieved by the DA algorithm is suboptimal in terms of social welfare, especially when there are many feasible stable matchings in the market. For instance, there has been a recent emergence of online matching markets, such as online labor markets (e.g., TaskRabbit, Upwork), online dating markets (e.g., Tinder, Match.com), and online crowdsourcing platforms (e.g., Amazon Mechanical Turk), where users do not know their preferences a priori and can repeatedly interact with the market to improve their matching quality [2]. The more a pair on both sides of the market gets matched, the more certain they become about their preferences. Therefore, an important question is how agents should interact with the market so that, in the absence of any coordination, they can learn their preferences as quickly as possible and achieve a stable matching through their collective behavior.
@@
In this work, we consider a two-sided matching market framed as a marriage problem, where the agents on one side are referred to as men and the agents on the other side as women. We assume that both men and women have preferences regarding the other side, and that women are aware of their ordinal preferences for men. However, men do not know their preferences for women and only learn them if they propose and get matched. In this case, the matched men receive a noisy estimate of their preferences. We assume that agents cannot observe any other information (e.g., who is rejected or accepted) and cannot coordinate in any way. The goal for men is to follow a decentralized and uncoordinated proposal strategy such that the entire market converges to a stable matching over time.
@@
One of the major challenges in learning a stable matching in markets with unknown preferences is handling collisions. More precisely, when multiple men propose to the same woman, only one of them gets matched and receives useful information, while all others are rejected and receive no information about their preferences for that woman. Therefore, resolving such collisions without coordination is a significant issue. Moreover, even if a man is matched with a woman, he receives a noisy utility drawn from an unknown distribution that may differ from his true preference. Consequently, a collision-avoidance process needs to be repeated many times before men can accurately estimate their true preferences. One of our goals in this work is to provide decentralized and uncoordinated algorithms that can effectively learn the true preferences while minimizing the number of collisions. Additionally, our work offers an alternative approach for measuring the closeness of the generated dynamics to any stable matching in general matching markets, rather than focusing on a specific stable matching in structured markets, as has been the main focus in past literature. To this end, we establish a connection between learning stable matchings in general markets with unknown preferences and learning Nash equilibria (NE) in noncooperative games. We leverage this connection to develop decentralized and uncoordinated algorithms for learning stable matchings in a principled manner.
I-A Related Work
@@
Stable matching was first introduced by [1] as a model for college admissions and to study the stability of marriages. Since then, there has been a tremendous effort to generalize and extend stable matchings to various market settings [3]. In their seminal work, Gale and Shapley provided a simple Deferred Acceptance (DA) algorithm in which men propose to their most preferred women, and the women reject all proposals except the one from their most preferred man. They showed that such a decentralized algorithm converges to a stable matching in at most $O(n^{2})$ steps, where $n$ is the total number of men and women in the market. Unfortunately, when the agents’ preferences are unknown, the DA algorithm is no longer guaranteed to converge to a stable matching without additional coordination. More precisely, if men fail to propose to their highest preferred women due to a lack of knowledge of their true preferences, their subsequent proposals may become unpredictable. For instance, if a man $m$ is rejected by a woman $w$ at a certain time $t$, it does not necessarily mean that $m$ should stop proposing to $w$. It could be that another man mistakenly proposed to $w$ at time $t$, causing $m$’s rejection. Therefore, to extend the DA algorithm to the case of unknown preferences, some form of coordination among the agents is required to avoid uninformative acceptance/rejection patterns. To address this issue, earlier works [4, 2, 5, 6] have proposed various coordination-based DA algorithms. The main idea is to develop a phase-dependent process where, in some phases, the agents primarily explore, and in other phases, they implement the DA algorithm based on current estimates of their unknown preferences.
@@
It is known that stable matchings with known preferences can be characterized using the extreme points of a fractional polytope that is totally dual integral [7, 8]. Moreover, stable matchings exhibit other useful properties that allow their geometry to be characterized using so-called rotations [9]. It was shown in [10] that uncoordinated random better-response dynamics converge to a stable matching with probability one. Subsequently, [11] provided an exponential lower bound for the worst-case convergence time of the uncoordinated random better/best response dynamics to a stable matching. However, all these results apply to the case when market preferences are fully known, allowing agents to compute their best or better responses and update their decisions accordingly.
@@
More recently, there has been a significant interest in matching markets with unknown preferences. However, depending on the information structure and the feedback received by the agents, one might expect a wide range of performance guarantees, which also depends on the type of algorithms followed by the agents (e.g., centralized/decentralized or coordinated/uncoordinated). The work in [12] devised a randomized polynomial-time centralized algorithm for finding a stable matching with unknown deterministic preferences. Each time, the algorithm proposes a matching and receives feedback in the form of a blocking pair, where an unmatched pair of man and woman $(m,w)$ is called blocking if both $m$ and $w$ prefer each other over their current partners. The study in [13] considers learning a specific stable matching with unknown stochastic preferences by adopting a suitable notion of stable regret, which measures the number of times that men propose to women other than their stable pair in that stable matching. The work in [2] addressed the problem of learning stable matchings in hierarchically structured markets where each submarket has a fixed pair (i.e., a pair of men and women who prefer each other the most in that submarket). They developed a phase-coordinated decentralized algorithm that achieves logarithmic regret in time with respect to the market’s unique stable matching. Additionally, [14] considered learning stable matchings in an uncoordinated and decentralized fashion. However, they require stronger assumptions on information feedback (e.g., a player observes the actions of other players in the previous round) and use a different performance metric than the one we consider in this work. We refer to [15] and [16] for other learning algorithms in two-sided matching markets with different performance guarantees. While all these works address the problem of learning a stable matching with unknown preferences, their algorithms differ substantially due to the information/feedback structure, the type of performance guarantee, and the level of coordination allowed among the agents.
I-B Contributions and Organization
@@
We consider the problem of learning stable matchings in matching markets where men do not know their preferences and aim to achieve a stable matching in a decentralized and uncoordinated manner. In particular, we focus on the weakest type of feedback that men can receive: a man observes a noisy version of his true preference only if his proposal is accepted, and receives no information otherwise. Our main objective is to provide a novel and principled approach to designing decentralized and uncoordinated learning algorithms for stable matchings by examining them through the lens of NE learning in noncooperative games. Our contributions can be summarized as follows:
  • (i)
@@
    We first provide a complete information continuous-action game formulation for the stable matching problem, namely the stable matching game, and show that its set of pure NE coincides with the set of stable matchings. Additionally, its mixed NE points can be rounded in a decentralized way to obtain a stable matching. This connection provides an alternative approach for measuring the closeness of the market dynamics to stable matchings through the lens of NE computation in games, thus extending the conventional notion of regret with respect to a specific stable matching in special markets to any stable matching in general markets.
  • (ii)
    Leveraging this game-theoretic formulation, we present a simple decentralized and uncoordinated algorithm, EXP, that globally converges to a stable matching in hierarchical markets. This algorithm achieves logarithmic regret in time, thereby extending the existing phase-coordinated algorithms in the literature to an uncoordinated one. The propose EXP algorithm can be viewed as the exponential weight learning algorithm adapted to the stable matching game. However, unlike the conventional analysis of EXP in the multi-arm bandit setting with a constant stepsize [17], the analysis of the EXP algorithm with time-varying parameters for the stable matching game is quite involved and requires a novel application of Freedman’s concentration inequalities for martingales in a hierarchical fashion.
  • (iii)
@@
    We then prove that the EXP algorithm always converges locally to a stable matching in general matching markets at an exponential rate. Moreover, we show that if the EXP algorithm converges globally with positive probability, it must converge to a stable matching.
  • (iv)
@@
    For general matching markets, it was previously unknown whether a fully decentralized and uncoordinated algorithm could globally learn a stable matching. Therefore, we complement our results by providing an alternative decentralized and uncoordinated algorithm that globally converges to a stable matching in general markets with arbitrarily high probability (albeit with a slower convergence rate), thanks to the weakly acyclic property of the stable matching game.
  • (v)
    Finally, in Appendix A, we address the problem of designing information feedback in matching markets and leverage our game-theoretic framework to identify conditions that facilitate the rapid global learning of stable matchings in general markets.
@@
The paper is organized as follows. In Section II, we formally introduce the problem. In Section III, we provide a complete information game-theoretic formulation for the stable matching problem and prove several characterization results for its set of NE points. In Section IV, we provide a decentralized and uncoordinated algorithm (EXP) that achieves logarithmic regret for hierarchical markets. In Section V, we show that EXP converges locally and exponentially fast to a stable matching in general matching markets. In Section VI, we provide an alternative decentralized and uncoordinated algorithm that globally converges to a stable matching in general matching markets with arbitrarily high probability. We conclude the paper by identifying some future research directions in Section VII. Omitted proofs and other supplementary materials are given in Appendix B.
II Problem Formulation
II-A Stable Matchings with Known Preferences
@@
Here, we first introduce the stable matching problem with known preferences [1]. In this problem, there are a set $M$ of men and a set $W$ of women, where by introducing dummy agents with appropriate preferences, without loss of generality we may assume $|M|=|W|=n$ [7] (we use the term “agents” to refer to either men or women). Each man $m\in M$ has a cardinal preference for each women $w\in W$, denoted by $\mu_{mw}\in(0,1]$, such if $\mu_{mw}>\mu_{mw^{\prime}}$, it means that $m$ prefers $w$ over $w^{\prime}$. Moreover, each woman $w$ has an ordinal preference over the men, and we write $m>_{w}m^{\prime}$ if woman $w$ strictly prefers $m$ over $m^{\prime}$. In this work we assume that no agent has ties in their preferences and we define
@@
II-B Stable Matchings with Unknown Preferences
@@
III A Game-Theoretic Formulation and Nash Equilibrium Characterization
In this section, we provide a complete information noncooperative game formulation for the stable matching problem with known preferences whose set of pure Nash equilibrium (NE) points coincides with the set of stable matchings. Later, we will show how to leverage this formulation to extend our results for learning stable matchings with unknown preferences. Such a formulation has three main advantages: i) it reduces the learning task in matching markets to learning NE in continuous-action concave games, ii) it captures the selfish behavior of men (players) and the feedback they receive through their payoff functions, and iii) it simplifies the combinatorial structure inherent in learning stable matchings to learning NE in continuous-action games. While analyzing the agents’ decisions directly on the discrete space of matchings seems complicated, which is why some earlier works need to restrict attention to special market structures [2], our game-theoretic reduction simplifies this task to some extent, thanks to the extensive literature on NE learning in continuous-action games.
III-A Stable Matching Game